OpenStudy (anonymous):

Could someone please check my answ? Will become fan!!! What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form. By using the y2-y1 over x2 - x1 the rise and run is 8 over 2 I still need help on writing the sope intercept.

OpenStudy (anonymous):

4 is the slope. So you are correct.

OpenStudy (anonymous):

Now you use one of your points to find the intercept

OpenStudy (anonymous):

(1, 4) 4=4(1)+b

OpenStudy (anonymous):

be looks to be 0

OpenStudy (anonymous):

So y=4x

OpenStudy (anonymous):

so y= 4x + b, what would (b) be ?

OpenStudy (anonymous):

draw the graph...u ll get the intercpt...

OpenStudy (anonymous):

See my previous posts. I go through that.

OpenStudy (anonymous):

y-y1=mx+c y-y1=4m(x-x1

OpenStudy (anonymous):

y=mx + b i think b = 0

OpenStudy (anonymous):

yes i have mentioned

OpenStudy (anonymous):

Yeah I just saw that thank you(: